Solve each linear equation. Show your work and check your answer.
Question #1 v-4(4-v) = -2 (2v-1)

Answer by stanbon(75887)

Question #1 v-4(4-v) = -2 (2v-1)
----------
v-16+4v = -4v+2
5v - 16 = -4v + 2
9v = 18
v = 2
===================
Cheers,
Stan H.